# On the Parallel Complexity of Identifying Groups and Quasigroups via Decompositions

### v1

@@ Abstract
  • We first consider the class $\mathcal{C}$ of groups that admit direct product decompositions, where each indecompsable factor is $O(1)$-generated, and either perfect or centerless. We show any group in $\mathcal{C}$ is identified by the $O(1)$-dimensional count-free Weisfeiler–Leman (WL) algorithm with $O(\log\log n)$ rounds, and the $O(1)$-dimensional counting WL algorithm with $O(1)$ rounds. Consequently, the isomorphism problem for $\mathcal{C}$ is in L. The previous upper bound for this class was $\textsf{TC}^{1}$, using $O(\log n)$ rounds of the $O(1)$-dimensional counting WL (Grochow and Levet, FCT 2023).
@@ 1 Introduction
The Quasigroup Isomorphism (QGpI) problem takes as input two finite quasigroups $G,H$ and asks if there exists an isomorphism $\varphi:G\to H$. When the quasigroups are given by their multiplication (Cayley) tables, it is known that QGpI belongs to $\textsf{NP}\cap\textsf{coAM}$. The generator-enumeration strategy has time complexity $n^{\log(n)+O(1)}$, where $n$ is the order of the group. In the setting of groups, this strategy was independently discovered by Felsch and Neubüser [FN70] and Tarjan (see [Mil78]). Miller subsequently extended this strategy to the setting of quasigroups [Mil78]. The parallel complexity of the generator-enumeration strategy has been gradually improved [LSZ77, Wol94, CTW13, Tan13]. Collins, Grochow, Levet, and Weiß recently improved the bound for the generator enumeration strategy to
@@
which can be simulated by depth-$4$ $\textsf{quasiAC}^{0}$ circuits of size $n^{O(\log^{4}n)}$ [CGLW24].
@@
In practice, such as working with computer algebra systems, the Cayley model is highly unrealistic. Instead, the groups are given by generators as permutations or matrices, or as black-boxes. In the setting of permutation groups, GpI belongs to NP [Luk93]. When the groups are given by generating sets of matrices, or as black-boxes, GpI belongs to $\textsf{Promise}\Sigma_{2}^{p}$ [BS84]; it remains open as to whether GpI belongs to NP or coNP in such succinct models. In the past several years, there have been significant advances on algorithms with worst-case guarantees on the serial runtime for special cases of this problem; we refer to [GQ17, DW22, GL23b] for a survey. To the best of our knowledge, polynomial-time isomorphism tests for quasigroups that are not groups have not been previously investigated.
@@
There is a natural divide-and-conquer strategy for GpI, which proceeds as follows. Given groups $G_{1},G_{2}$, we first seek to decompose the groups in terms of normal subgroups $N_{i}\trianglelefteq G_{i}$ ($i\in[2]$) and the corresponding quotients $G_{i}/N_{i}$. It is then necessary (but not sufficient) to decide if $N_{1}\cong N_{2}$ and $G_{1}/N_{1}\cong G_{2}/N_{2}$. This strategy has led to algorithmic advances for groups admitting nice decompositions, including for instance, direct product decompositions [KN09, Wil12, BS22, GL23b], coprime [Gal09, QST11, BQ12, GL23b] and tame [GQ15] extensions, and central extensions [LW12, BMW17, IQ19, BGL+19, GQ17]. In case of quasigroups the situation is made more complicated by the lack of normal substructures; one has to factor by congruences instead. (Recall that a congruence of an algebra $A$ is an equivalence relation on $A$ that is compatible with all operations of $A$.) We therefore focus on the isomorphism problem for quasigroups that admit representations in terms of more structured algebras, such as groups and their automorphisms.
@@
We first investigate the complexity of computing direct product decompositions. Brachter and Schweitzer [BS22] showed that Weisfeiler–Leman (WL) can identify direct products in polynomial-time, provided it can also identify the indecomposable direct factors in polynomial-time. Grochow and Levet [GL23b] parallelized this result, essentially showing that WL can detect the indecomposable direct factors using only $O(\log n)$ rounds. On the other hand, Grohcow and Levet (ibid.) showed that the analogous result fails to hold for the count-free variant of WL. While cyclic groups are identified by the count-free $2$-WL, there exist infinite families $(G_{m},H_{m})_{m\in\mathbb{N}}$ of non-isomorphic Abelian groups that require $\Omega(\log|G_{m}|)$-dimensional (and hence, $\Theta(\log n)$-dimensional) count-free WL to distinguish between $G_{m}$ and $H_{m}$.
It is thus natural to inquire as to the ability of count-free WL to identify groups that admit nice direct product decompositions. The Remak–Krull–Schmidt theorem provides that for any two fully-refined direct product decompositions $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ and $\mathcal{T}=\{T_{1},\ldots,T_{k}\}$ of a group $G$, there exists a permutation $\pi\in\text{Sym}(k)$ such that $S_{i}\cong T_{\pi(i)}$ for all $i\in[k]$. In general, however, $S_{i}$ and $T_{\pi(i)}$ need not be the same subgroup (setwise). On the other hand, it is well-known that when $S_{i}$ is centerless, it appears (setwise) in every fully-refined direct product decomposition of $G$. The same holds true when $S_{i}$ is perfect (see Corollary 3.6). For this reason, we investigate the count-free Weisfeiler–Leman dimension for the following class of groups.
@@
For $k\geq 2,r\geq 0$, let $(k,r)$-WL denote running $k$-dimensional Weisfeiler–Leman for $r$ rounds. Our first main result is the following.
@@
Grochow and Levet previously established that the counting $(O(1),O(\log n))$-WL Version II identifies the groups in $\mathcal{C}$. In light of the parallel WL implementation due to Grohe and Verbitsky [GV06] (see also Section 2.5), Grochow and Levet obtained that isomorphism testing between a group $G\in\mathcal{C}$ and an arbitrary group $H$ is in $\textsf{TC}^{1}$. Theorem 1.2(a) already improves this bound to $\textsf{AC}^{1}$.
@@
We also compare Theorem 1.2(a) to the work of Collins and Levet [CL24], who showed that count-free WL can be fruitfully leveraged as a subroutine for isomorphism testing of direct products of non-Abelian simple groups. Note that such groups all belong to $\mathcal{C}$. Collins and Levet showed that the count-free $(O(1),O(\log\log n))$-WL Version I will distinguish between two non-isomorphic direct factors. They then used $O(\log n)$ co-nondeterministic bits and a single Majority gate to verify the multiplicities of the isomorphism type for each direct factor appeared the same number of times in both groups. This yields a bound of $\forall^{\log n}\textsf{MAC}^{0}(\textsf{FOLL})$ (following the notation of [CGLW24]). Theorem 1.2(a) shows that count-free WL on its own can identify this class of groups. On the other hand, the bound obtained by Collins and Levet of $\forall^{\log n}\textsf{MAC}^{0}(\textsf{FOLL})$ is stronger than the bound of $\textsf{AC}^{1}$ afforded by Theorem 1.2(a). This brings us to our second main result.
@@
The count-free $(O(1),O(\log\log n))$-WL Version I algorithm identifies every group $G$, that decomposes as a direct product of non-Abelian simple groups.
@@
Our next result concerns the canonization problem. For a class $\mathcal{K}$ of objects, a canonical form is a function $F:\mathcal{K}\to\mathcal{K}$, such that for all $X,Y\in\mathcal{C}$, we have (i) $X\cong F(X)$, and (ii) $X\cong Y\iff F(X)=F(Y).$ Isomorphism testing reduces to canonization, and the converse remains open. Here, we establish the following:
@@
Let $d>0$ be a constant. There exists a uniform $\textsf{AC}^{3}$ algorithm such that the following holds. Let $G$ be a group of order $n$ that admits a fully-refined direct product decomposition, such that each indecomposable direct factor is $d$-generated. Our algorithm computes a canonical labeling of $G$.
@@
Central quasigroups play an important role since they are precisely the Abelian objects (in the sense of universal algebra) in the variety of quasigroups [Sze98]. See [KN, KN71, Dra09, Smi76, Smi90] for an introduction to central quasigroups and [Hou12, Kir95, SV15] for enumeration of certain classes of central quasigroups up to isomorphism.
@@
In contrast, the complexity of isomorphism testing for Abelian groups has been well-investigated [LSZ77, Vik96, Sav80, Kav07, Ili85, Vil94, CTW13, GL23b, CGLW24], resulting in a linear time serial algorithm [Kav07] and parallel bounds of $\forall^{\log\log n}\textsf{MAC}^{0}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$, the latter of which is a proper subclass of $\mathsf{L}$ [CGLW24].
@@ 1.1 Further Related Work
There is a long history of work on developing efficient parallel (NC) algorithms for special cases of Graph Isomorphism. We refer to [LRS24] for a survey. In comparison, the work on NC algorithms for (Quasi)Group Isomorphism is comparatively nascent. Much of the work revolves around parallelizing the generator-enumeration strategy (ibid.), culminating in bounds of L for $O(1)$-generated groups [Tan13] and $\textsf{SAC}^{1}$ for $O(1)$-generated quasigroups [Wag10]. Prior to this paper (Theorem 1.10), no other special family of quasigroups was known to even admit an polynomial-time isomorphism test. Other families of groups known to admit NC isomorphism tests include Abelian groups [Vil94, CTW13, GL23b, CGLW24], graphical groups arising from the CFI graphs [BS20, CL24, Col23], coprime extensions $H\ltimes N$ where $H$ is $O(1)$-generated and $N$ is Abelian [GL23b] (parallelizing a result from [QST11]), groups of almost all orders [CGLW24] (parallelizing [DW22]), and Fitting-free groups [GJL25] (parallelizing [BCQ12]).
@@
The use of combinatorial techinques for GpI is relatively new; and to the best of our knowledge, was initiated in [LQ17, BGL+19]. Brachter and Schweitzer [BS20] subsequently initiated the study of Weisfeiler–Leman applied directly to the Cayley table. In particular, they established that graphical groups [Mek81, HQ21] arising from the CFI graphs [CFI92] have WL-dimension at most $3$. There has been considerable work building on the framework introduced by Brachter and Schweitzer– see [BS22, Bra23, GL23a, GL23b, CL24, Col23, CRP24, Vag23, GJL25]. It remains open whether the class of finite groups has WL-dimension $o(\log n)$.
@@ 2.1 Groups and Quasigroups
Let $G$ be a group. Given $g,h\in G$, the commutator $g,h]:=ghg^{-1}h^{-1}$. For sets $X,Y\subseteq G$, the commutator subgroup $X,Y]:=\langle\{[g,h]:g\in X,h\in Y\}\rangle$. We say that $G$ is perfect if $G=[G,G]$, and that $G$ is centerless if $Z(G)=1$. Let $d(G)$ be the minimum size taken over all generating sets of $G$. A basis of an Abelian group $A$ is a set of generators $\{a_{1},\ldots,a_{k}\}$ such that $A=\langle a_{1}\rangle\times\cdots\times\langle a_{k}\rangle.$

### v2

@@ Abstract
  • We first consider the class of groups that admit direct product decompositions, where each indecompsable factor is $O(1)$-generated, and either perfect or centerless. We show any group in this class is identified by the $O(1)$-dimensional count-free Weisfeiler–Leman (WL) algorithm with $O(\log\log n)$ rounds, and the $O(1)$-dimensional counting WL algorithm with $O(1)$ rounds. Consequently, the isomorphism problem for this class is in L. This improves upon the previous upper bound of $\textsf{TC}^{1}$, which was obtained using $O(\log n)$ rounds of the $O(1)$-dimensional counting WL (Grochow and Levet; FCT 2023, J. Comput. Syst. Sci. 2026).
@@ 1 Introduction
The Quasigroup Isomorphism (QGpI) problem takes as input two finite quasigroups $G,H$ and asks if there exists an isomorphism $\varphi:G\to H$. When the quasigroups are given by their multiplication (Cayley) tables, it is known that QGpI belongs to $\textsf{NP}\cap\textsf{coAM}$ (see e.g., [GMW86]). The generator-enumeration strategy has time complexity $n^{\log(n)+O(1)}$, where $n$ is the order of the group. In the setting of groups, this strategy was independently discovered by Felsch and Neubüser [FN70] and Tarjan (see [Mil78]). Miller subsequently extended this strategy to the setting of quasigroups [Mil78]. The parallel complexity of the generator-enumeration strategy has been gradually improved [LSZ77, Wol94, CTW13, Wag10, Tan13]. Collins, Grochow, Levet, and Weiß recently improved the bound for the generator enumeration strategy to
@@
which can be simulated by depth-$4$ $\textsf{quasiAC}^{0}$ circuits of size $n^{O(\log n)}$ [CGLW25].
@@
In practice, such as working with computer algebra systems, the Cayley model is highly unrealistic. Instead, the groups are given by generators as permutations or matrices, or as black-boxes. In the setting of permutation groups, GpI belongs to NP [Luk93]. When the groups are given by generating sets of matrices, or as black-boxes, GpI belongs to $\textsf{Promise}\Sigma_{2}^{p}$ [BS84]; it remains open as to whether GpI belongs to NP or coNP in such succinct models. In the past several years, there have been significant advances on algorithms with worst-case guarantees on the serial runtime for special cases of this problem; we refer to [GQ17, DW22, GL26] for a survey. To the best of our knowledge, polynomial-time isomorphism tests for quasigroups that are not groups have not been previously investigated.
@@
There is a natural divide-and-conquer strategy for GpI, which proceeds as follows. Given groups $G_{1},G_{2}$, we first seek to decompose the groups in terms of normal subgroups $N_{i}\trianglelefteq G_{i}$ ($i\in[2]$) and the corresponding quotients $G_{i}/N_{i}$. It is then necessary (but not sufficient) to decide if $N_{1}\cong N_{2}$ and $G_{1}/N_{1}\cong G_{2}/N_{2}$. This strategy has led to algorithmic advances for groups admitting nice decompositions, including for instance, direct product decompositions [KN09, Wil12, BS22, GL26], coprime [Gal09, QST11, BQ12, GL26] and tame [GQ15] extensions, and central extensions [LW12, BMW17, IQ19, BGL+19, GQ17]. In case of quasigroups the situation is made more complicated by the lack of normal substructures; one has to factor by congruences instead. (Recall that a congruence of an algebra $A$ is an equivalence relation on $A$ that is compatible with all operations of $A$.) We therefore focus on the isomorphism problem for quasigroups that admit representations in terms of more structured algebras, such as groups and their automorphisms.
@@
We first investigate the power of the Weisfeiler–Leman (WL) algorithm to identify groups based on their direct product decompositions. Weisfeiler–Leman is a key combinatorial algorithm used in the setting of Graph Isomorphism. Brachter and Schweitzer recently provided three adaptations to the setting of groups, specified by their multiplication (Cayley) tables [BS20]. In this work, we will use only the first two adaptations, which Brachter and Schweitzer refer to as WL Versions I and II. We will provide a brief overview of WL here, and we refer to Section 2.3 for a precise formulation. For $k\geq 2$, the $k$-dimensional Weisfeiler–Leman algorithm iteratively colors $k$-tuples of group elements in an isomorphism-invariant manner. At each refinement step, each $k$-tuple receives a new color based on (i) its current color, and (ii) the multiset of colors from the “nearby” $k$-tuples. Additionally, we will be interested in the weaker count-free variant of Weisfeiler–Leman, in which the refinement steps only take into account the set of colors from the “nearby” $k$-tuples, rather than the full multiset. Grohe and Verbitsky [GV06] observed that WL can be effectively parallelized (see Section 2.5 for further discussion). Consequently, controlling both the dimension and number of iterations can yield improvements in the parallel complexity of isomorphism testing.
We now turn to discussing direct product decompositions. Kayal and Nezhmetdinov [KN09] and Wilson [Wil12] previously exhibited polynomial-time algorithms to compute direct product decompositions of finite groups into indecomposable direct factors (we refer to such decompositions as fully refined– see Section 3.1 for more background on direct products). For group isomorphism testing, the main benefit of these results is that, for the purposes of getting polynomial-time isomorphism tests, they reduce the problem to the case of directly indecomposable groups.
Brachter and Schweitzer [BS22] established a WL-analogue, by relating the Weisfeiler–Leman dimension for an arbitrary group $G$, to that of $G$’s indecomposable direct factors. Grochow and Levet [GL26] parallelized this result, essentially showing that WL can detect the indecomposable direct factors using only $O(\log n)$ rounds. While WL does not explicitly return a direct product decomposition itself, for the purposes of isomorphism testing, WL achieves essentially the same benefit of reducing to the directly indecomposable case– in $\textsf{TC}^{1}$ (at the cost of increasing the dimension by $1$ and adding $O(\log n)$ rounds) [GL26].
On the other hand, Grohcow and Levet (ibid.) showed that the analogous result fails to hold for the count-free variant of WL. While cyclic groups are identified by the count-free $2$-WL, there exist infinite families $(G_{m},H_{m})_{m\in\mathbb{N}}$ of non-isomorphic Abelian groups that require $\Omega(\log|G_{m}|)$-dimensional (and hence, $\Theta(\log|G_{m}|)$-dimensional) count-free WL to distinguish between $G_{m}$ and $H_{m}$.
Motivated by this gap, we investigate the ability of count-free WL to identify groups that admit nice direct product decompositions. The Remak–Krull–Schmidt theorem provides that for any two fully-refined direct product decompositions $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ and $\mathcal{T}=\{T_{1},\ldots,T_{k}\}$ of a group $G$, there exists a permutation $\pi\in\text{Sym}(k)$ such that $S_{i}\cong T_{\pi(i)}$ for all $i\in[k]$. In general, however, $S_{i}$ and $T_{\pi(i)}$ need not be the same subgroup (setwise). On the other hand, it is well-known that when $S_{i}$ is centerless, it appears (setwise) in every fully-refined direct product decomposition of $G$. The same holds true when $S_{i}$ is perfect (see Corollary 3.6). For this reason, we investigate the distinguishing power of the count-free Weisfeiler–Leman algorithm for the following class of groups.
@@
As controlling the WL-dimension and the number of rounds can yield improvements in the parallel complexity for isomorphism testing [GV06] (see Section 2.5 for more discussion), we will use the following notation. Let $k\geq 2,r\geq 0$, and let $(k,r)$-WL denote running $k$-dimensional Weisfeiler–Leman for $r$ rounds.
Our first main result is the following.
@@
We refer to Section 2.3 for a precise formulation of WL Versions I and II.
Grochow and Levet previously established that the counting $(O(1),O(\log n))$-WL Version II identifies the groups in $\mathcal{C}$. In light of the parallel WL implementation due to Grohe and Verbitsky [GV06] (see also Section 2.5), Grochow and Levet obtained that isomorphism testing between a group $G\in\mathcal{C}$ and an arbitrary group $H$ is in $\textsf{TC}^{1}$, which improved upon the previous bound of P. Theorem 1.2(a) already improves this bound from $\textsf{TC}^{1}$ to $\textsf{AC}^{1}$.
@@
We also compare Theorem 1.2(a) to the work of Collins and Levet [CL24], who showed that count-free WL can be fruitfully leveraged as a subroutine for isomorphism testing of direct products of non-Abelian simple groups. Note that such groups all belong to $\mathcal{C}$. Collins and Levet showed that the count-free $(O(1),O(\log\log n))$-WL Version I will distinguish between two non-isomorphic direct factors. They then used $O(\log n)$ universally quantified co-nondeterministic bits and a single Majority gate to verify the multiplicities of the isomorphism type for each direct factor appeared the same number of times in both groups. Following the notation of [CGLW25], we refer to this corresponding complexity class as $\forall^{\log n}\textsf{MAC}^{0}(\textsf{FOLL})$. Theorem 1.2(a) shows that count-free WL on its own can identify this class of groups. On the other hand, the bound obtained by Collins and Levet of $\forall^{\log n}\textsf{MAC}^{0}(\textsf{FOLL})$ is stronger than the bound of $\textsf{AC}^{1}$ afforded by Theorem 1.2(a). This brings us to our second main result.
@@
The count-free $(O(1),O(\log\log n))$-WL Version I algorithm identifies every group $G$ that decomposes as a direct product of non-Abelian simple groups.
@@
As $\mathsf{FOLL}$ cannot compute Parity or Majority [FSS84], we have that $\mathsf{FOLL}\subsetneq\forall^{\log n}\textsf{MAC}^{0}(\textsf{FOLL})\subseteq\textsf{AC}^{1}$.
Our next result concerns the canonization problem. For a class $\mathcal{K}$ of objects, a canonical form is a function $F:\mathcal{K}\to\mathcal{K}$, such that for all $X,Y\in\mathcal{K}$, we have (i) $X\cong F(X)$, and (ii) $X\cong Y\iff F(X)=F(Y).$ Isomorphism testing reduces to canonization, and the converse remains open.
We will, in particular, design an algorithm that computes canonical labelings. Precisely, if $X,Y\in\mathcal{K}$ have order $n$, our algorithm will return labelings $\lambda_{X}:X\to[n]$ and $\lambda_{Y}:Y\to[n]$ such that $X\cong Y$ if and only if $\lambda_{Y}^{-1}\circ\lambda_{X}$ is an isomorphism between $X$ and $Y$. We establish the following:
@@
Let $d>0$ be a constant. There exists a uniform $\textsf{AC}^{3}$ algorithm such that the following holds. Let $G$ be a group of order $n$ that admits a fully-refined direct product decomposition such that each indecomposable direct factor is $d$-generated. Our algorithm computes a canonical labeling of $G$.
@@
Central quasigroups play an important role since they are precisely the Abelian objects (in the sense of universal algebra) in the variety of quasigroups [Sze98]. See [KN71a, KN71b, Dra09, Smi76, Smi90] for an introduction to central quasigroups and [Hou12, Kir95, SV15] for enumeration of certain classes of central quasigroups up to isomorphism.
@@
In contrast, the complexity of isomorphism testing for Abelian groups has been well-investigated [LSZ77, Vik96, Sav80, Kav07, Ili85, Vil94, CTW13, GL26, CGLW25], resulting in a linear time serial algorithm [Kav07] and parallel bounds of $\forall^{\log\log n}\textsf{MAC}^{0}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$, the latter of which is a proper subclass of $\mathsf{L}$ [CGLW25].
@@ 1.1 Further Related Work
There is a long history of work on developing efficient parallel (NC) algorithms for special cases of Graph Isomorphism. We refer to [LRS24] for a survey. In comparison, the work on NC algorithms for (Quasi)Group Isomorphism is comparatively nascent. Much of the work revolves around parallelizing the generator-enumeration strategy (ibid.), culminating in bounds of L for $O(1)$-generated groups [Tan13] and $\textsf{SAC}^{1}$ for $O(1)$-generated quasigroups [Wag10]. Prior to this paper (Theorem 1.10), no other special family of quasigroups (beyond certain families of groups) was known to even admit an polynomial-time isomorphism test. Other families of groups known to admit NC isomorphism tests include Abelian groups [Vil94, CTW13, GL26, CGLW25], graphical groups arising from the CFI graphs [BS20, CL24, Col23], coprime extensions $H\ltimes N$ where $H$ is $O(1)$-generated and $N$ is Abelian [GL26] (parallelizing a result from [QST11]), groups of almost all orders [CGLW25] (parallelizing [DW22]), and Fitting-free groups [GJL25] (parallelizing [BCQ12]).
@@
The use of combinatorial techinques for GpI is relatively new; and to the best of our knowledge, was initiated in [LQ17, BGL+19]. Brachter and Schweitzer [BS20] subsequently initiated the study of Weisfeiler–Leman applied directly to the Cayley table. In particular, they established that graphical groups [Mek81, HQ21] arising from the CFI graphs [CFI92] have WL-dimension at most $3$. There has been considerable work building on the framework introduced by Brachter and Schweitzer– see [BS22, Bra23, GL25, GL26, CL24, Col23, CRP24, Vag23, GJL25]. It remains open whether the class of finite groups has WL-dimension $o(\log n)$.
@@ 2.1 Groups and Quasigroups
Let $G$ be a group. Let $Z(G)$ denote the center of $G$. Given $g,h\in G$, the commutator $g,h]:=ghg^{-1}h^{-1}$. For sets $X,Y\subseteq G$, the commutator subgroup $X,Y]:=\langle\{[g,h]:g\in X,h\in Y\}\rangle$. We say that $G$ is perfect if $G=[G,G]$, and that $G$ is centerless if $Z(G)=1$. Let $d(G)$ be the minimum size taken over all generating sets of $G$. Fix $d\in\mathbb{N}$; we say that a class $\mathcal{K}$ of groups is $d$-generated if for each group $G\in\mathcal{K}$, $d(G)\leq d$. A basis of an Abelian group $A$ is a set of generators $\{a_{1},\ldots,a_{k}\}$ such that $A=\langle a_{1}\rangle\times\cdots\times\langle a_{k}\rangle.$ We say that a subgroup $N\leq G$ is characteristic if for all $\alpha\in\operatorname{Aut}(G)$, $\alpha(N)=N$.
For a group $G$ and a subset $S\subseteq G$, with $1\not\in S$ and $S=S^{-1}$, the Cayley graph $\text{Cay}(G,S)$ has vertex set $G$. There is an edge $\{g,h\}$ in $\text{Cay}(G,S)$ if and only if there exists some $s\in S$ such that $gs=h$.